OpenStudy (carissa15):

I have a precipitation question. I have started through it but keep getting the incorrect answer. Please help?

OpenStudy (carissa15):

\[ksp =(2x)^2 *x =7.1*10^-7\] and then got the answer x=1.525*10^-4

OpenStudy (kkutie7):

it might help if you were to write out the chemical equation first.

OpenStudy (anonymous):

can i help

OpenStudy (vincent-lyon.fr):

In the equation, what concentrations are known are what concentrations are unknown?